Consider the following. (a) Find the force Fx associated with the potential-energy function U Ax5, where A is a constant.(Use the following as necessary: A and x.) (b) At what value(s) of x does the force equal zero? XE
Answer #1

Given that

the potential energy U=Ax^5

(a) ans

now we find the force Fx

the force Fx=du/dx

=d/dx(Ax^5)

=5Ax^4

(b) ans

the displacement X=0 m

the force F(x)=Ax^4

force F(0)=A(0)^5=0